The nurse is to infuse 1000mL 0.9 NS at 125 mL/hr using macrodrip (15 gtt/mL) and microdrip (60 gtt/mL) IV tubing. Calculate the drip rates for both.
Answer #1

1.Drip rate = Total amount of fluid * drop factor / total number of hours * 60

= 125*15 / 1*60

= 1875 / 60

= 31.25

= 31 drops/minute.

We have to keep 31 drops /minute in macrodrip set to infuse 125ml per hour.

2. Drip rate = Total amount of fluid * Drip factor / total number of hours *60

= 125 * 60 / 1 *60

= 7500 / 60

= 125 drops per minute

We have to keep 125 drops per minute in microdrip set to infuse 125ml per hour.